The model y = 135 + 25 x represent the cost of renting a car, where x is the number of days the car is rented.

a- What is the y-intercept for this model? Describe what this number represents in terms of the variables in the model.

b- What is the slope for this model? Describe does this number mean in terms of renting a car.

c- Evaluate the model for x=4. Describe what the ordered pair in terms of renting a car.


Answer by usyim88hk(157)

You can see the slope and the y-intercept from the equation y = 135+25x easily because it is in the slope-intercept form already...
:
y=mx+b
:
where m is the slope and b is the y-intercept
:
So the y-intercept is 135 and it means that you must pay $135 first in order to rent a car
:
The slope is 25 and it means that for each day the driver has to pay $25
:
If x is 4
Then y = 135+25(4)
y = 135+100
y = 235
So this means that if the number of days the driver rent the car is 4, then he will have to pay $235